# VoicePilot: Harnessing LLMs as Speech Interfaces for Physically Assistive Robots

### v1

@@ Abstract.
Physically assistive robots present an opportunity to significantly increase the well-being and independence of individuals with motor impairments or other forms of disability who are unable to complete activities of daily living. Speech interfaces, especially ones that utilize Large Language Models (LLMs), can enable individuals to effectively and naturally communicate high-level commands and nuanced preferences to robots. Frameworks for integrating LLMs as interfaces to robots for high level task planning and code generation have been proposed, but fail to incorporate human-centric considerations which are essential while developing assistive interfaces. In this work, we present a framework for incorporating LLMs as speech interfaces for physically assistive robots, constructed iteratively with 3 stages of testing involving a feeding robot, culminating in an evaluation with 11 older adults at an independent living facility. We use both quantitative and qualitative data from the final study to validate our framework and additionally provide design guidelines for using LLMs as speech interfaces for assistive robots. Videos and supporting files are located on our project website [^1] https://sites.google.com/andrew.cmu.edu/voicepilot/
conference: User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA
Figure 1. Top: Our work introduces an iteratively constructed framework for integrating LLMs as speech interfaces for assistive robots. Middle: We develop a speech interface using the framework and evaluate it with 11 older adult participants at an independent living facility. Bottom: Using insights from the study, we present a reflection and 5 design guidelines.
@@
Figure 1. Top: Our work introduces an iteratively constructed framework for integrating LLMs as speech interfaces for assistive robots. Middle: We develop a speech interface using the framework and evaluate it with 11 older adult participants at an independent living facility. Bottom: Using insights from the study, we present a reflection and 5 design guidelines.
@@ 1. Introduction
Physically assistive robots, such as the adaptive eating robot Obi (Obi, [n. d.]), shown in Fig. 2, can enable individuals with impairments to perform a range of self-care and household tasks (Ciocarlie et al., 2012; Hawkins et al., 2014; King et al., 2012; Nanavati et al., 2023b). Assistive interfaces can allow individuals with varying forms of disability to control such robots. Interfaces often have physical requirements, dictating which individuals can effectively use them. For example, individuals without the ability to move their head may use eye tracking with a web-based interface (Chen et al., 2013; Cabrera et al., 2021) while those with head motion may prefer inertial interfaces that capture their head movements and directly map them to robot motion (Padmanabha et al., 2023, 2024).
@@
Among these interfaces, speech excels as an option for those with the ability to speak, as it allows individuals to naturally provide robots with both high-level commands and nuanced preferences or customizations. Speech interfaces for assistive robots have been explored extensively for control of both mobile robots and robot arms (Pulikottil et al., 2018; Lauretti et al., 2017; House et al., 2009) and can be more intuitive than existing interface options like joysticks (Poirier et al., 2019). With the advent of Large Language Models (LLMs) and ongoing research combining LLMs with robots, robust speech interfaces for robots are becoming increasingly viable in the near future. Existing works have presented frameworks and guidelines for integrating LLMs with robots for high-level task planning and code generation (Singh et al., 2023; Liang et al., 2023; Vemprala et al., 2023a). However, these works focus on improving prompt engineering and code generation and lack guidelines regarding the human-centric considerations necessary when utilizing LLMs as assistive interfaces.
@@
To address this gap, we develop a framework and design guidelines for integrating LLMs as speech interfaces for physically assistive robots. Our presented framework underwent extensive iterative development. First, we leveraged insight from existing literature to develop an initial version. We applied this Version 1 framework to integrate an existing LLM, GPT-3.5 Turbo from OpenAI (gpt, [n. d.]), with the commercial feeding robot Obi (Obi, [n. d.]). After testing this interface with lab members, we refined the framework, iterated on our LLM-based speech interface for the Obi, and demonstrated it to community members at a disability awareness event (Version 2). We used insights from the demonstration to develop a final framework and interface, evaluated through a user study involving 11 older adults at an independent living facility (Version 3). Lastly, we use both quantitative and qualitative data from the study with older adults to present design guidelines.
@@
Figure 3. Our final framework consisting of 9 components is shown. The color and annotation at the bottom right of each numbered component indicates in which framework iteration it was added.
Our final framework consisting of 9 components is shown. The color and annotation at the bottom right of each numbered component indicates in which framework iteration it was added. The first 5 components: Environment Description, Robot Functions, Function Applications, Code Specifications, and Safety were added in Version 1. Robot Variables was added in Version 2. Lastly, Instructional Materials, User Control Functions, and Feedback were added in Version 3.
Figure 3. Our final framework consisting of 9 components is shown. The color and annotation at the bottom right of each numbered component indicates in which framework iteration it was added.
@@ 2.1. LLMs for Robotics
There have been numerous advancements in using LLMs for robotics (Firoozi et al., 2023; Kawaharazuka et al., 2024; Zhang et al., 2023; Vemprala et al., 2023b). For the scope of this paper, we will focus on works that use off-the-shelf LLMs for high-level task planning through code generation, as these works are most pertinent to using LLMs as assistive interfaces. First, Liang et al. use codewriting LLMs in conjunction with perception and control APIs to generate both reactive and waypoint based trajectories for various robotic platforms (Liang et al., 2023). Vemprala et al. present design principles and a pipeline for integrating ChatGPT with various robots (quadcopter and 6 degree-of-freedom arm) and robotic simulation environments (Vemprala et al., 2023a). Other work has proposed using ChatGPT as a user interface for a daVinci surgical robot and quadrupeds (Pandya, 2023; Macdonald et al., 2024). Similarly, Singh et al. present design and prompting guidelines and evaluate their methodology using various LLMs in simulation and with robot arms for 70 household tasks (Singh et al., 2023). Lastly, Arenas et al. present PromptBook, a set of 7 guidelines for having LLMs generate code for robots that combines instruction-based and example-based prompting, chain-of-thought, interleaved states, and human feedback (Arenas et al., 2023). Many of these past works blur the line between using LLMs to generate code for robots and using LLMs as an interface for users to communicate with and command a robot. While there is some overlap between the two, prompting LLMs for high level planning and code generation is just one aspect of designing a speech interface. Our paper focuses on human-centric considerations of using LLMs as speech interfaces, such as safety and customizability, which are crucial for a positive user experience, especially in assistive contexts.
@@
Past works have also explored using LLMs with social and service robots (Wu et al., 2023; Hu et al., 2024; Kim et al., 2024; Kian et al., 2024). Wu et al. (Wu et al., 2023) focus on learning personal preferences through language input for TidyBot, a cleaning robot, while Hu et al. (Hu et al., 2024) benchmark LLM performance for code generation for various service tasks. Both papers focus on how LLMs can be utilized to perform a specific task instead of the interface design and the human-robot interaction. Lastly, Kim et al. explore using LLMs with a social robot and focus on conversational tasks, such as having the robot guide the participant in making a drink (Kim et al., 2024). The authors compare their robot against text and voice based agents and present design insights from a thematic analysis of the data obtained during the human study. Most of these insights are only applicable to social, conversational robots and do not generalize to physically assistive robots and tasks such as robot-assisted feeding. In comparison to all past work on using LLMs for robots, we iterate on our presented framework through multiple rounds of testing with users and we introduce design guidelines derived from our final human study.
2.2. Speech Interfaces for Robot-Assisted Feeding
@@ 3.1. Framework Development
Our Version 1 framework identifies five preliminary components, shown and described in Fig. 3. We additionally describe how these preliminary components were motivated from past literature:
@@
  • Robot Functions: Past papers on using LLMs for high-level task planning and code generation suggest defining and listing a set of robot functions in the prompt so that the LLM can access these function names and use them to control the robot (Vemprala et al., 2023a; Arenas et al., 2023; Singh et al., 2023). Similar robot functions are required for LLM-based assistive speech interfaces. For example, for an assistive cleaning robot, you may include a function to perceive objects and a function to grasp an object which can be used in conjunction by an LLM to generate code to detect and pick up toys from the floor of a cluttered room.
@@ 3.2. Implementation
We addressed the components as follows: to implement the Environment Description, within our prompt, we included details on the task at hand, feeding the user, and the working environment, such as the foods in each of the bowls. We incorporated the Robot Functions and Function Applications components by providing the names of three high-level robotic control functions in the prompt, along with a short description of what each function did and how it should be used. These three functions moved the robot arm along hard-coded trajectories and were described as follows: obi.scoop_from_bowlno(bowlno) moved the robotic arm to the specified bowl and scoops food from it; obi.move_to_mouth() moved the robot to the user’s mouth position; and obi.scrape_down_bowlno(bowlno) moved the robot to the specified bowl and subsequently scraped food from the sides to the center of that bowl so that the robot could pick up additional food. At the end of the prompt, to satisfy the Code Specifications component, we included sentences instructing GPT to provide Python code for each request. Lastly, to address the Safety component, each function only moved the robot along a predefined trajectory, as opposed to allowing GPT to directly dictate the joint angles of Obi’s arm. This prevents GPT from generating unsafe or unanticipated trajectories. Choosing to use the Obi robot, which is specifically designed to be safe for individuals with motor impairments, instead of other robotic feeding systems, is another implementation decision for enhancing safety.
@@ 3.3. Pilot Testing with Lab Members
We conducted pilot testing of our Version 1 interface for the Obi robot within our research team and members of our lab group.
First, four lab members, separate from the immediate research team, were presented with short videos depicting the robot performing four tasks which were (1) feeding a single scoop from one bowl, (2) feeding multiple scoops from one bowl in succession, (3) feeding scoops from multiple, different bowls in succession, and (4) scraping down the sides of a bowl to move food to the center of the bowl. They were then asked to each provide three different ways of phrasing commands to the robot to complete each task, intending to gather a broader set of phrasings that was not influenced by the researcher’s vocabulary; this resulted in a total of 48 phrasings. Lastly, we gauged all individuals on improvements and features they would like to see implemented in the system.
We tested the accuracy of the LLM using the 48 collected command phrasings. We recorded which phrasings did not elicit the intended code from the LLM, and we subsequently analyzed these mistakes to determine if there were common patterns. Next, we made minor modifications to the prompt, tackling one mistake at a time. Subsequently, we tested the 48 commands with the new prompt and compared the LLM’s new output to the output from the previous prompt to assess whether our modification rectified the issue and whether it introduced any unforeseen side effects. We repeated this process of testing, analysis, and modification multiple times until all of the mistakes were addressed.
@@ 4.1. Framework Iteration
Based on our pilot testing, we noted instances where GPT made errors, enabling us to identify gaps or areas within the framework that required iteration. First, the Code Specifications in our prompt were not sufficiently detailed, as GPT would frequently attempt to import modules that had already been imported or did not exist, attempt to call functions that did not exist, define unnecessary new functions to carry out the user’s request, or simply not generate code. Thus, we modified the description of Code Specifications in the framework to express the importance of explicitly outlining how GPT should write code.
We additionally noticed that GPT would sometimes call robot functions incorrectly, primarily in the following three ways. First, if a user requested a scoop of food, GPT would frequently scoop from a bowl without bringing the food up to the user’s mouth. Additionally, if a user said “scrape down the yogurt bowl and then feed me some”, GPT would frequently scrape down the sides of the bowl and then immediately move to the user’s mouth, without scooping to actually pick up a bite. Lastly, if a user asked to be fed multiple times in succession, GPT would frequently not pause long enough at the user’s mouth for the user to comfortably take the bite, instead moving immediately from the user’s mouth to the next bowl to scoop from. For these reasons, we expanded the Function Applications component of the framework to emphasize the need to not only describe what each of the functions do, but to describe how the functions relate to each other and to provide examples of how the functions should be combined. To further alleviate these issues, we modified the Robot Functions component, to suggest the merging of functions that are always called in sequence. These solutions generalize to other assistive robots: for example, a mobile manipulator in the home may need to do multiple actions in sequence, such as grabbing a cup and filling it with water before bringing it to the user. For this case, if GPT frequently skipped the filling step, specifying the logical order of functions in the prompt or even merging the “grabbing cup” and “filling water” functions would lead to better performance.
@@
Due to the user safety implications of the new Robot Variables component which gives the LLM speech interface control of motion variables such as speed, we proactively modified the Safety component of the framework to suggest restriction of robot variables such as speed to conservative ranges.
@@ 4.2. Implementation
To adhere to the updated Function Applications component, we added more information to the prompt about how the robot functions should be combined. Specifically, we placed additional emphasis on Obi moving to the user’s mouth if the user asked to be fed or have a bite. Additionally, we specifically instructed GPT to pause between bites if users asked to be fed multiple times in succession to give users the chance to eat each bite; without this pause, GPT would immediately move from the user’s mouth to the next scoop, combining the functions incorrectly as the user would not be able to be fed properly. We set the default delay to be 4 seconds and users could customize this delay if desired.
Additionally, to address the altered Robot Functions component, the function to scrape down the bowl was modified to always scoop from the same bowl immediately after to enhance GPT’s accuracy. This change was incorporated as we realized that whenever Obi was requested to scrape down a bowl, it was with the intention of scooping food immediately from that bowl after. Previously, GPT often failed to comprehend this sequence, leading to instances where it would incorrectly scrape down the bowl’s sides and then move directly to the user’s mouth without scooping, despite the our intention for Obi to scoop as well.
Lastly, to incorporate the new Robot Variables component, we defined three new global variables that affect the way in which Obi carries out its movements. These include obi.speed, which sets the speed at which Obi moves and ranges from 0 degrees per second (slowest) to 80 (fastest); obi.accel, which determines the rate at which Obi accelerates and ranges from 0 degrees per second squared (slowest) to 250 (fastest); and obi.deep_scoop, which dictates the size of the scoops Obi takes with bigger scoops if true and smaller scoops if false. The speed ranges and scoop depths were chosen with the modified Safety component in mind; specifically, in terms of speed, we found that moving the robot at full speed (120 degrees per second) presents potential risks, such as food spillage. Lastly, we added the names of each of these robot variables to the prompt, along with descriptions of what they affect, their ranges, and their default values.

### v2

@@ Abstract.
Physically assistive robots present an opportunity to significantly increase the well-being and independence of individuals with motor impairments or other forms of disability who are unable to complete activities of daily living. Speech interfaces, especially ones that utilize Large Language Models (LLMs), can enable individuals to effectively and naturally communicate high-level commands and nuanced preferences to robots. Frameworks for integrating LLMs as interfaces to robots for high level task planning and code generation have been proposed, but fail to incorporate human-centric considerations which are essential while developing assistive interfaces. In this work, we present a framework for incorporating LLMs as speech interfaces for physically assistive robots, constructed iteratively with 3 stages of testing involving a feeding robot, culminating in an evaluation with 11 older adults at an independent living facility. We use both quantitative and qualitative data from the final study to validate our framework and additionally provide design guidelines for using LLMs as speech interfaces for assistive robots. Videos, code, and supporting files are located on our project website [^1] https://sites.google.com/andrew.cmu.edu/voicepilot/
assistive robotics, large language models (LLMs), speech interfaces
journalyear: 2024copyright: rightsretainedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676401isbn: 979-8-4007-0628-8/24/10ccs: Computer systems organization External interfaces for roboticsccs: Computing methodologies Speech recognition
Figure 1. Top: Our work introduces a framework that was iteratively developed for integrating LLMs as speech interfaces for assistive robots. Middle: We develop a speech interface for a feeding robot using the framework and evaluate it with 11 older adults at an independent living facility. Bottom: Using insights from the study, we present a reflection and 5 design guidelines.
[…]
@@ 1. Introduction
Physically assistive robots, such as the adaptive eating robot Obi (Obi, [n. d.]), shown in Fig. 2, can enable individuals with impairments to perform a range of self-care and household tasks (Ciocarlie et al., 2012; Hawkins et al., 2014; King et al., 2012; Nanavati et al., 2023b; Yang et al., 2023). Assistive interfaces can allow individuals with varying forms of disability to control such robots. Interfaces often have physical requirements, dictating which individuals can effectively use them. For example, individuals without the ability to move their head may use eye tracking with a web-based interface (Chen et al., 2013; Cabrera et al., 2021) while those with head motion may prefer inertial interfaces that capture their head movements and directly map them to robot motion (Padmanabha et al., 2023, 2024).
@@
Among these interfaces, speech excels as an option for those with the ability to speak, as it allows individuals to naturally provide robots with both high-level commands and nuanced preferences or customizations. Speech interfaces for assistive robots have been explored extensively for control of both mobile robots and robot arms (Pulikottil et al., 2018; Lauretti et al., 2017; House et al., 2009) and can be more intuitive than existing interface options like joysticks (Poirier et al., 2019). With the advent of Large Language Models (LLMs) and ongoing research combining LLMs with robots, robust speech interfaces for robots are becoming increasingly viable in the near future. Existing works have presented frameworks and guidelines for integrating LLMs with robots for high-level task planning and code generation (Singh et al., 2023; Liang et al., 2023; Vemprala et al., 2023a; Brohan et al., 2023; Zhang et al., 2023; Zhao et al., 2023; Ahn et al., 2024; Wang et al., 2024; Mahadevan et al., 2024). These works concentrate on prompt engineering and code generation to enhance the accuracy of large language models (LLMs). No prior works involve human subjects interacting directly with an LLM-integrated robot for physically assistive tasks. Consequently, they lack insight and guidelines on the human-centric considerations essential for using LLMs as assistive interfaces.
@@
To address this gap, we develop a framework and design guidelines for integrating LLMs as speech interfaces for physically assistive robots. Our presented framework underwent extensive iterative development and consists of components related to prompt engineering and system rollout, the process of deploying the LLM-based speech interface and robot with users. First, we leveraged insight from existing literature to develop an initial version. We applied this Version 1 framework to integrate an existing LLM, GPT-3.5 Turbo from OpenAI (gpt, [n. d.]), with the commercial feeding robot Obi (Obi, [n. d.]). After testing this interface with lab members, we refined the framework, iterated on our LLM-based speech interface for the Obi, and demonstrated it to community members at a disability awareness event (Version 2). We used insights from the demonstration to develop a final framework and interface, evaluated through a user study involving 11 older adults at an independent living facility (Version 3). Lastly, we use both quantitative and qualitative data from the study with older adults to present design guidelines.
@@
Figure 3. Our final framework consisting of 9 components is shown. The color and annotation at the top left of each numbered component indicates whether the component is related to prompt engineering (PE), system rollout (SR), or both prompt engineering and system rollout (PS). System rollout refers to the process of deploying the LLM-based speech interface and robot with users. The annotation at the bottom right of each numbered component shows which framework iteration it was added in.
[…]
Our final framework consisting of 9 components is shown. The color and annotation at the top left of each numbered component indicates whether the component is related to prompt engineering (PE), system rollout (SR), or both prompt engineering and system rollout (PS). The color and annotation at the bottom right of each numbered component indicates in which framework iteration it was added. The first 5 components: Environment Description, Robot Functions, Function Applications, Code Specifications, and Safety were added in Version 1. Robot Variables was added in Version 2. Lastly, Instructional Materials, User Control Functions, and Feedback were added in Version 3.
@@ 2.1. LLMs for Robotics
There have been numerous advancements in using LLMs for robotics (Firoozi et al., 2023; Kawaharazuka et al., 2024; Vemprala et al., 2023b; Singh et al., 2023; Liang et al., 2023; Vemprala et al., 2023a; Brohan et al., 2023; Zhang et al., 2023; Zhao et al., 2023; Ahn et al., 2024; Wang et al., 2024; Mahadevan et al., 2024). For the scope of this paper, we will focus on works that use off-the-shelf LLMs for high-level task planning through code generation, as these works are most pertinent to using LLMs as assistive interfaces. Liang et al. utilize codewriting LLMs with perception and control APIs to create reactive and waypoint-based trajectories for various robotic platforms, while Vemprala et al. develop design guidelines and a pipeline for integrating ChatGPT with different robots and simulation environments (Liang et al., 2023; Vemprala et al., 2023a). Simarily, Singh et al. and Arenas et al. provide prompting guidelines and methodologies for using LLMs in robotic tasks (Singh et al., 2023; Arenas et al., 2023). Other work has used ChatGPT as an interface for surgical robots and quadrupeds (Pandya, 2023; Macdonald et al., 2024). These prior works primarily focus on high level planning and code generation which is just one aspect of designing a speech interface. Most utilize prompt engineering and evaluate their LLMs using accuracy, specifically if the LLM succeeds in generating the “correct” code that the researcher expects based on the provided prompt and command. While some papers include human studies to evaluate their methods, none evaluate their system with participants who interacted directly with the LLM-integrated robot. In comparison, our work includes multiple rounds of direct interaction between the robot and participants including with older adults who are close to the target population of individuals who need feeding assistance. This process allows us to identify human-centric elements essential to positive user perception of the system.
@@
Past works have also explored using LLMs with social and service robots (Wu et al., 2023; Hu et al., 2024; Kim et al., 2024; Kian et al., 2024). Wu et al. (Wu et al., 2023) focus on learning personal preferences through language input for TidyBot, a cleaning robot, while Hu et al. (Hu et al., 2024) benchmark LLM performance for code generation for various service tasks. Both papers focus on how LLMs can be utilized to perform a specific task and similar to other papers, do not have participants interacting directly with the robot. The absence of these human studies limits the insights into improving the general interface design and human-robot interaction. Lastly, Kim et al. explore using LLMs with a social robot and focus on conversational tasks, such as having the robot guide the participant in making a drink (Kim et al., 2024). The authors compare their robot against text and voice based agents and present design insights from a thematic analysis of the data obtained during the human study. Most of their findings are only applicable to social, conversational robots and do not generalize to physically assistive robots and tasks such as robot-assisted feeding. In comparison to all past work on using LLMs for robots, we iterate on our presented framework through multiple rounds of testing with users and we introduce design guidelines derived from our final human study.
2.2. Robot-Assisted Feeding
@@ 3.1. Framework Development
Our Version 1 framework identifies five preliminary components, shown and described in Fig. 3. We describe how these preliminary components were motivated from past literature:
@@
  • Robot Functions: Past papers on using LLMs for high-level task planning and code generation suggest defining and listing a set of robot functions, named the Robot API, in the prompt so that the LLM can access these function names and use them to control the robot (Vemprala et al., 2023a; Arenas et al., 2023; Singh et al., 2023). Similar robot functions are required for LLM-based assistive speech interfaces. For example, for an assistive cleaning robot, you may include a function to perceive objects and a function to grasp an object which can be used in conjunction by an LLM to generate code to detect and pick up toys from the floor of a cluttered room.
@@ 3.2. Implementation
We addressed the components as follows: to implement the Environment Description, within our prompt, we included details on the task at hand, feeding the user, and the working environment, such as the foods in each of the bowls, as the Obi robot does not have a built-in perception system. We incorporated the Robot Functions and Function Applications components by providing the names of three high-level robotic control functions in the prompt, along with a short description of what each function did and how it should be used. These three functions moved the robot arm along hard-coded trajectories and were described as follows: obi.scoop_from_bowlno(bowlno) moved the robotic arm to the specified bowl and scooped food from it; obi.move_to_mouth() moved the robot to the user’s mouth position; and obi.scrape_down_bowlno(bowlno) moved the robot to the specified bowl and subsequently scraped food from the sides to the center of that bowl so that the robot could pick up additional food. We designed the Robot API to include these three functions because they correspond to actions performed by the commercially available Obi robot. Specifically, the robot’s default two-button interface allows users to select the bowl with one button and scoop/feed themselves with the other. Additionally, the scraping motion occurs automatically after every few scoops.
At the end of the prompt, to satisfy the Code Specifications component, we included sentences instructing GPT to provide Python code for each request. Lastly, to address the Safety component, each function only moved the robot along a predefined trajectory, as opposed to allowing GPT to directly dictate the joint angles of Obi’s arm. This prevents GPT from generating unsafe or unanticipated trajectories. Choosing to use the Obi robot, which is specifically designed to be safe for individuals with motor impairments, instead of other robotic feeding systems, is another implementation decision for enhancing safety.
@@ 3.3. Pilot Testing with Lab Members
We conducted pilot testing of our Version 1 interface for the Obi robot within our research team and members of our lab group. Four lab members watched videos of the robot performing four feeding tasks and suggested different command phrasings to accomplish the task, resulting in 48 phrases. These commands were tested for accuracy, with mistakes analyzed and prompts adjusted iteratively. Later, three researchers directly interacted with the robot to further assess and improve the system, focusing on user comfort during assisted feeding. More details on this process is included in Appendix A.1.
@@ 4.1. Framework Iteration
Based on our pilot testing, we noted instances where GPT made errors, enabling us to identify gaps or areas within the framework that required iteration. Specifically, the Code Specifications were not detailed enough, causing GPT to import nonexistent modules and define unnecessary new functions. To address this, the framework was modified to emphasize the importance of explicitly outlining how GPT should write code. Additionally, the Function Applications component was expanded to better describe the relationships between functions and provide examples, while the Robot Functions component was modified to suggest merging common sequential functions to improve performance in tasks that required multiple robot actions. More information on these changes can be found in Appendix A.2.
@@
Due to the user safety implications of the new Robot Variables component which gives the LLM speech interface control of motion variables such as speed, we proactively modified the Safety component of the framework to suggest restriction of robot variables such as speed to conservative ranges and to check the validity of robot variables in LLM-outputted code to ensure they are within expected bounds.
@@ 4.2. Implementation
In summary, we updated the prompt to emphasize the production of responses exclusively in Python code, without importing extra modules or creating new functions, as specified in the revised Code Specifications component. To adhere to the updated Function Applications component, we added detailed instructions on how the robot functions should be combined, emphasizing Obi’s movement to the user’s mouth when asked to be fed, and included a 4-second default pause between bites for better user experience, with customization options for the delay. Additionally, the Robot Functions component was modified to ensure that after scraping down a bowl, Obi would immediately scoop from the same bowl to improve accuracy, addressing previous issues where the sequence was misunderstood. For the Robot Variables component, we added new variables for Obi’s movements (speed, acceleration, and scoop size) to enhance customizability, with post-processing by the system to ensure safety. More information on these changes can be found in Appendix A.3.